OpenStudy (anonymous):

9x^2=10 I need to see the steps so I get it

OpenStudy (anonymous):

\[9x ^{2}=10\]

OpenStudy (saifoo.khan):

[Math Processing Error]

OpenStudy (diyadiya):

If you're solving for x \[\Huge {9x^2=10 \\x^2 = \frac{10}{9} }\]\[\Huge{ \sqrt{x^2}=\sqrt{\frac{10}{9}}\\ x=\frac{\sqrt{10}}{\sqrt{9}}= \frac {\sqrt{10}}{3}}\]